# Are Biological Systems More Intelligent Than Artificial Intelligence?

### v6

@@ 7 Hybrid Creation as Weak Boundary-Condition Design
Consider the aforementioned example of the organoid adapted to play the game Pong with notable sample efficiency relative to conventional reinforcement-learning baselines [53]. Hybrid systems like organoids are not fixed interpreters of software-level policy, but collectives whose lower-level parts retain bottom-up agency. In such systems, hybrid creation is the act of designing an interface and environment that reshapes the feasible policy space of the biological components. Rather than “programming” the parts directly, the designer engineers boundary conditions so that the remaining viable collective policies implement a target input-output relation at the interface. Practically, hybrid creation is iterative:
@@ 8 Conclusions and Future Research
Biological systems often adapt efficiently because they delegate adaptive control across scales. Using Stack Theory [19], I formalised delegation as the depth to which adaptation is permitted in a stack of abstraction layers and proved The Law of the Stack (Theorem 5.1). Utility one layer up is bounded by the weakness of the realised policy one layer down. Delegation is a structural prerequisite for high-level adaptability.
@@
This constraint is compatible with active inference. Under a simple viability prior, minimising variational free energy is equivalent to maximising entropy over viable futures (w-maxing; ESM “Viability, lived history, and variational free energy”). Corollary 5.1 shows that free energy minimisation inherits the same bottleneck: abstract-layer free energy cannot be reduced without sufficient lower-layer slack.
@@
The implication for AI is not that computation cannot be intelligent, but that contemporary AI stacks tend to concentrate adaptation at high software layers while freezing lower layers, thereby throttling achievable weakness and adaptation efficiency. For hybrid agents (organoids, soft robots, and human–AI organisations) the design problem becomes boundary-condition engineering: constrain parts just enough to elicit target collective behaviour while preserving a non-empty intersection of viable policies (collective identity). When that intersection collapses, the system can fragment into competing subagents (Section 6).
@@
Future work should test these predictions by varying which layers are permitted to adapt (e.g. hardware-in-the-loop learning, morphologically adaptive substrates, and biologically integrated controllers) and measuring resulting degrees of freedom, free energy, and robustness. More broadly, this paper is a stepping stone toward a unified account of delegation, weakness/max-entropy heuristics, and multiscale agency in the larger Stack Theory programme [19].
A practical design implication is that, when adaptability is desired, constraints should be weakened as far as possible without breaking the system’s intended function. Constraints that are not strictly necessary relative to the design goal should be removed (What counts as “strictly necessary” is relative to design goals and value judgements. See ESM Supplementary Note S4.). Humans often prefer ‘additive’ solutions [1], meaning if a problem can be solved by adding a rule or subtracting a rule, we tend to add a rule. However The Law of the Stack suggests a preference for subtractive solutions might be beneficial in many circumstances.
@@
To speculate about future research beyond experimental verification, Stack Theory is a lens that can be applied generally. For example, a number of software-first approaches to AI already approximate aspects of this delegation picture, including federated learning [52], constructivist AGI programmes [90, 42], and active inference [39, 40]. An expanded discussion (including AERA, Hyperon/ActPC-Chem, and links to scale-free active inference) is provided in the ESM (Supplementary Note S6). Stack Theory might also inform the design of ‘soft robots’ [58]. Soft robots are a proposed class of robots made of ‘soft’ materials that adapt. I have formalised adaptive materials as adaptation lower in the stack, which can be used as a formal foundation for soft robotics. Self-assembling nanoparticle networks [24] are a step in this direction. They can be used to form reconfigurable logic gates [70], facilitating adaptation at very low levels of abstraction. Taken to its logical conclusion, future AI may more closely resemble artificial life [85, 45, 46]. Finally, Stack Theory may be extended to economic and organisational systems. For example, free-market trade could be seen as a classic argument for delegation. As Hayek pointed out [44], a great deal of relevant information is spread among a population, unavailable to a central planner. This motivates delegation of control, to shorten feedback loops. However, the present formalism treats resources implicitly via task correctness, so connecting these arguments to economic dynamics would require extending Stack Theory to explicit time-series [23] and resource accumulation. Additional speculative extensions and applications (including organisational and economic systems, AI safety regulation, and temporal extensions of stacks) are collected in ESM Note S9.
In summary, I have applied Stack Theory to explore how delegation of control affects adaptability. The Law of the Stack shows how adaptability at higher levels of abstraction is constrained by adaptability at lower levels, and delegation is structurally necessary to minimise free energy. This lays a foundation for a broad range of future theoretical and empirical research, showing how systems of any kind might be made more robust and adaptable.
@@ Acknowledgements
First, I’d like to thank the anonymous peer reviewers for their detailed and helpful feedback. Second, I’d like to thank Anna Ciaunica, Peter Watts, Sean Welsh, Ricard Solé, Karl Friston, Antonio Damasio, Kingson Man, Paolo Milani, Ben Goertzel, Kristinn R. Thórisson, Elija Perrier and Noel Hinton for their support and feedback on this paper.
S1. Working definition of intelligence
There are many definitions of intelligence, and the merits of each are beyond the scope of this paper. Mine is based on [19], which frames intelligence as the ability to complete a wide range of tasks, which is sample and energy efficiency in adaptation. It is similar to Wang’s definition of intelligence as adaptation with limited resources [96], and Hutter’s earlier definition of intelligence as ‘the ability to satisfy goals in a wide range of environments’. Mainstream AI often operationalizes ‘intelligence’ as general-purpose task performance. By that standard, one may argue that current large language models are ‘more intelligent’ than many biological systems. However that definition is fundamentally flawed, as it fails to express the comparative efficiency with which humans acquire general purpose competence [16, 96, 26]. In theory even a lookup table or rote-learning response can possess general purpose competence, given enough resources. This paper instead treats intelligence as the cost of adaptation (sample and energy efficiency), and the results should be read accordingly.
S2. Levels of abstraction beyond computing
The notion of levels of abstraction is more typically used in pure computer science settings, but can be applied to any system [19]. A higher level of abstraction is the behaviour of the layers below. For example, an organ like a heart is a behaviour of cells, human is a behaviour of some organs, and language is a behaviour of humans.
S3. Weakness and w-maxing (extended intuition)
Weakness is formally defined in the appendix. Related work proved w-maxing to be optimal regardless of the abstraction layer [18, 15, 16, 19]. In narrative terms, assume two policies $x$ and $y$ both constrain the system to examples of “correct” behaviour given past inputs, but differ in weakness. If $x$ is weaker than $y$, that means it commits to less, but achieves the same end. For example, if I assume the lawyer was paid at least $\mathdollar 5$, I am more weakly constrained than if I assume that lawyer was paid at least $\mathdollar 6$.
This example also helps separate weakness from simplicity. The difference between $\mathdollar 5$ and $\mathdollar 6$ is not primarily about how hard the rule is to write down. It is about how many further commitments remain compatible with the rule, given the agent’s available distinctions. Weakness is defined relative to an abstraction layer. The layer fixes what counts as a meaningful refinement because it fixes what the system can jointly discriminate and jointly realise. If the embodiment changes, then the induced vocabulary changes, and so does the completion structure that weakness counts.
Many familiar heuristics can be seen as special cases of the same idea once this completion structure is made explicit. If one treats all completions that remain possible at a layer as equally plausible, then choosing a correct policy that maximises $|\mathrm{Ext}(\pi)|$ is equivalent to maximising entropy over viable futures inside that layer. This parallels least commitment and version space intuitions, but it is stated directly in terms of the agent’s embodied language rather than in terms of an encoding chosen by an observer. That is what lets weakness be used across layers in Stack Theory, where a policy at one layer induces the next layer vocabulary via the abstractor map.
In the language of agency, weakness can be read as a quantitative proxy for autonomy at a layer. A weaker policy leaves more degrees of freedom (more “ways of being right”) while still satisfying the same observed constraints. In multiscale and hybrid systems this autonomy matters because it increases the chance that component policies remain compatible with a higher-level collective policy, preserving collective identity.
As an epistemological razor, this means “explanations should be no more specific than necessary” or alternatively “rules should constrain no more than necessary” (see Bennett’s Razor in [15]). This is a bit like the principle of maximum entropy [50, 49], difference being that w-maxing also accounts for the abstraction layer, which is important because abstraction can render complexity subjective [17].
For a rigorous definition of learning and efficiency, see definitions in the online repository for the thesis [21]. It includes proofs showing w-maxing is both necessary and sufficient to maximise the probability of ‘generalising’ from one task to another, and thus adaptability.
S4. Normativity and “strict necessity” in design
As Hume argued, a description of what ought to be cannot be derived from only descriptions of what is. “Strictly necessary” is a value judgement, or an implicit choice of what ought to be. In the thesis [19] on which this is based this ought from which other descriptions of ought follow is derived from the state transitions of the environment. These transitions preserve some aspects of that environment, while destroying others. Here however I am more concerned with the design of systems which serve a human purpose. Strictly necessary is a matter of desired functionality.
S5. Conceptual motivation for the formal setting
Stack Theory frames the description of an objective reality in terms of choices of subjective lens (an abstraction layer) [19]. The full justification for these definitions is out of scope and addressed at length elsewhere [16, 19, 18]. However I will provide a brief overview. To make claims that hold regardless of the abstraction layer, we must rely only on what holds for every abstraction layer. From our subjective perspective we can’t know what is at the bottom of our stack (it may be infinite). I call this bottom layer the environment. Whatever the environment is, there must be change or difference for there to be anything at higher levels of abstraction. Hence there must be a set of states where each state is one point of difference (Nothing is assumed about what these states might ontologically be or epistemologically contain.). Because these states must be mutually exclusive, this frames time as difference. Each possible sequence of states is a timeline or possible world. Truth is defined ontologically, relative to states. A set of states is effectively a set of truth conditions, so a possible fact is called a declarative program, and a program is true given a state $\phi$ if it contains that state. Every aspect of the environment is just a set of programs. Everything that exists is a set of facts, and each timeline preserves different sets of facts. A timeline is like an opinion. As time progresses some aspects of the environment are preserved and others are destroyed. Each timeline preserves different aspects, expressing a different “ought” which related work uses to explain the origins of goal directed behaviour [19]. If the set of states is infinite then there are infinite possible worlds. Assume I have a robotic system. I get an abstraction layer for the system by choosing a finite aspect $\mathfrak{v}$ of programs called a vocabulary. Everything the abstraction layer can express is a subset of its vocabulary. For example if a system has an arm it can raise or lower, then there are programs in $\mathfrak{v}$ which correspond to actuation of the arm in some form, though it need not be neat symbolic representations like ’raise arm’ or ’lower arm’. Since states are mutually exclusive, $\mathfrak{v}$ gives us an embodied formal language. I do not mean like the words humans use to communicate. I mean a language of embodiment, where every statement in that language constrains rather than describes the system, like the aforementioned boundary conditions on slime mold [62] (Using this frame cognition can be formalised in terms that align with both enactivist [86] and pancomputationalist [73] perspectives. Hence, cognition within Stack Theory is also called Pancomputational Enactivism[16].)
S6. Extended related work on delegated software architectures and Active Inference/FEP
There are examples of systems that take something like the present approach, albeit within the confines of software abstraction layers.
Federated machine learning delegates control to some extent (e.g. shifting data ownership and parts of optimisation to edge devices), which can improve privacy and efficiency [52]. Constructivist theories of artificial general intelligence emphasise multilayered self-organisation at the level of software while acknowledging at least some of the constraints imposed by hardware [90, 42]. The Auto-catalytic Endogenous Reflective Architecture (AERA) is a goal-driven ampliative reasoning system [64] based on constructivist principles that relies on causal knowledge representation to achieve autonomy [89], “growing” knowledge in a collective meta-model through cumulative continuous learning [88], and explicitly targeting adaptability [87, 91, 33, 80] (See also http://www.openaera.org.). ActPC-Chem [43] is a recent theory behind the distributed Hyperon architecture [41] that was influenced by theoretical and methodological insights from AERA.
Active Inference/FEP in relation to MCL.
Active Inference (and the Free Energy Principle, FEP) provides a normative account of perception and action in which an agent maintains a generative model and selects actions to minimise variational free energy (and, by extension, expected free energy) [39, 40, 71]. The present paper addresses a different axis. It states a structural constraint on stacks of abstraction layers, namely that what can be achieved at layer i+1i{+}1 (utility $\epsilon(\gamma_{i+1})$) is bounded by the extension (weakness) of the realised policy at layer $i$ (Theorem 5.1). In this sense, MCL is about where adaptation is implemented (and how much latitude lower layers retain), rather than about the particular inferential objective used within a layer. An Active Inference agent implemented on a predominantly fixed computational substrate can still be bottlenecked by static lower layers. Conversely, a delegated multiscale substrate can in principle host a range of learning rules, including Active Inference.
Potential complementarity and open questions.
Several lines of work suggest points of contact, without implying formal equivalence. ActPC-Chem frames delegation in explicitly Active Inference terms [43, 39, 40]. “Scale-free” active inference proposes that similar dynamical principles repeat across scales [36, 37], which is conceptually close to the MCA picture and to the idea that policy alignment must be maintained across layers to preserve collective identity. Delegated [68, 69, 67] or federated [38] Active Inference approaches may therefore be promising candidates for building efficient hybrid systems at scale. Separately, work on “mortal computation” seeks to approach thermodynamic limits (e.g. the Landauer bound) by reducing memory and processor separation and associated bottlenecks [66, 5, 97], which resonates with the present paper’s emphasis on efficiency and with the intuition that pushing adaptation downward can reduce overhead. However, the theories still differ in their primitives (e.g. tasks and weakness versus generative models and free energy). Even so, a simple mapping is available when viability is treated as a hard constraint and preferences are uniform across viable futures, see the Mathematical Results subsection “Viability, lived history, and variational free energy”. A deeper identification in the general non-uniform case remains open.
S7. Mission command example (Nordbat 2)
A well documented example of the effectiveness of mission command in facilitating adaptation is Nordbat 2, a joint peacekeeping force deployed to Bosnia in the early 90s. Mission command allowed individual units to deviate from convention, adapt to local conditions and adopt a more ‘trigger-happy’ disposition that proved to be substantially more effective than the more rigid, rule based approach taken by the Dutch forces operating in the same region [47, 48].
Strategic-level intent and a small set of constraints are communicated downward, while tactical-level decision-makers fill in the remaining degrees of freedom on the ground. In this reading, the bridge between leadership and soldiers is not a detailed plan but a constraint system that keeps local improvisation aligned with higher-level purpose. The purpose of this note is not to treat Nordbat 2 as empirical evidence for Theorem 5.1, but to provide an intuition-building picture of what Stack Theory means by delegation [47, 48].
Mapped to the formalism, the strategic layer specifies a parent task $\omega$ (intent, mandate, and coordination requirements), while the tactical layer faces a family of child tasks $\alpha\sqsubset\omega$ that vary with local information and contingencies. Mission command enlarges the set of admissible completions by avoiding over-specification of the realised policy. It preserves a larger extension $|\mathrm{Ext}(\pi_{i})|$ at the lower layer, enabling more distinct higher-layer behaviours to be expressed one layer up (Theorem 5.1).
For hybrid and cyberphysical design, the analogy is that effective “control” can often be achieved by engineering boundary conditions (interfaces, reward signals, resource limits, safety constraints) rather than by prescribing micro-actions. Over-prescriptive constraints risk collapsing the intersection of viable component policies (loss of collective identity), whereas intent-plus-boundaries can preserve slack for local adaptation while maintaining coherence.
S8. Candidate empirical and simulation tests
The main text uses analogies (mission command, economic delegation, and distributed swarms) as intuition-building thought experiments rather than as evidence. It is primarily conceptual. This note lists candidate tests that target the core claim of Theorem 5.1. The goal is to compare systems under controlled variation of where adaptation is permitted in the stack.
1. Agent-based simulations with tunable delegation.
Construct an agent-based model of a collective. Include a delegation parameter that controls whether decisions are taken locally or routed through higher-level coordination. Expose the collective to abrupt environmental shifts. Measure sample efficiency, energy use, robustness, and identity preservation. Test whether deeper delegation improves efficiency in the sense predicted by reduced weakness of lower-layer realised policies.
2. Comparative benchmarking across stack depth.
Compare learning systems that differ only in where adaptation is permitted. Hold the top-layer objective fixed and vary whether lower layers are static or reconfigurable. For example, compare fixed-substrate deployment against deployment that allows adaptation of firmware, routing, or morphology. Measure adaptation speed and compute the effective induced vocabulary size at each layer when possible. This isolates the stack-level delegation hypothesis from algorithm choice at the top.
3. Hybrid boundary-condition sweeps.
For hybrid agents such as simulated mixtures of human and AI agents or engineered organoids, treat alignment as hybrid agent design. Sweep boundary conditions that prune low-level policy spaces. Quantify the trade-off between task performance at the interface and preservation of non-empty collective policy intersection. Look for phase transitions where the intersection collapses, which corresponds to loss of collective identity.
4. Interaction with Active Inference/FEP.
Implement Active Inference style control at one or more layers. Run the same comparison in an architecture that differs only in whether lower layers can adapt. This disentangles gains due to the inferential objective from gains due to deeper delegation.
S9. Extended applications and speculative directions
This note collects material that is peripheral to the main theorem but may be useful for future-work framing. These applications are intended as design hypotheses rather than as evidence for Theorem 5.1.
Organisational and economic systems.
Free-market trade is a great example of how delegation can facilitate adaptability in a large collective. As Hayek pointed out [44], not all of the information is available to a central planner and so control is best delegated to the individual. However, the accumulation of resources within my formalism is implicit, rather than explicit. I have been concerned only with whether a system satisfies whatever arbitrary notion of correct I choose (i.e. does it achieve homeostatic and reproductive goals in whatever amount of time I care about?), rather than how it evolves over time as a dynamical system. In other words, I have concerned myself only with whether a system arrives on time at the destination I want, not the journey it took to get there. Future research might extend Stack Theory to formalise time-series, or the accumulation of various resources. For example, a stack could represent a sequence of tasks over time where higher layers are later in time. Earlier I mentioned that task correctness can denote more holistic notions of emergent features robust to lower-level variation. A time-series model could describe the conditions under which such invariants emerge. It could be used to explore how the delegation of control, distribution of work and the accrual of resources change over time. For example under some conditions a free market may degrade into a maladaptive, top-down system. Those parts deprived of control or resources may then not be able to construct a policy that aligns with the larger whole, and so splinter off. To preserve the integrity of the system, policies which might support the desired invariants could be modelled and evaluated. Policy could be located along axes of distribution and delegation, like an alternative political compass. In a simple zero-sum game like a race, one example of such a policy is adaptive rubber banding [61]. A game mechanic (e.g. in Mario Kart) where success accrues risk as well as reward, keeping the game competitive.
Regulation and over-constraint risks.
The cancer analogy suggests a general caution. Guardrails that over-constrain a system may precipitate the failure mode they aim to prevent by collapsing the intersection of viable policies. For example, survey of 3,306 Germans found that restrictions promoting carbon-neutral lifestyles would trigger strong negative responses because they ‘restrict freedom’ [77]. This may be most relevant to proposals for safeguarded AGI and related governance approaches [29, 28, 3, 42], given the formal results pertain specifically to a generalisation-optimal model of intelligence [18]. These remarks are speculative and are intended as design hypotheses rather than claims about current regulatory regimes.

### v7

@@ 7 Hybrid Creation as Weak Boundary-Condition Design
Consider the aforementioned example of the organoid adapted to play the game Pong with notable sample efficiency relative to conventional reinforcement-learning baselines [khajehnejad2024]. Hybrid systems like organoids are not fixed interpreters of software-level policy, but collectives whose lower-level parts retain bottom-up agency. In such systems, hybrid creation is the act of designing an interface and environment that reshapes the feasible policy space of the biological components. Rather than “programming” the parts directly, the designer engineers boundary conditions so that the remaining viable collective policies implement a target input-output relation at the interface. Practically, hybrid creation is iterative:
@@ 8 Conclusions and Future Research
Biological systems often adapt efficiently because they delegate adaptive control across scales. Using Stack Theory [bennett2025thesis], I formalised delegation as the depth to which adaptation is permitted in a stack of abstraction layers and proved The Law of the Stack (Theorem 5.1). Utility one layer up is bounded by the weakness of the realised policy one layer down. Delegation is a structural prerequisite for high-level adaptability.
@@
This constraint is compatible with active inference. Under a simple viability prior, minimising variational free energy is equivalent to maximising entropy over viable futures (w-maxing; see ESM “Viability, lived history, and variational free energy”). Corollary 5.1 shows that free energy minimisation inherits the same bottleneck, which is that abstract-layer free energy cannot be reduced without sufficient lower-layer slack.
@@
The implication for AI is not that computation cannot be intelligent, but that contemporary AI stacks tend to concentrate adaptation at high software layers while freezing lower layers, thereby throttling achievable weakness and adaptation efficiency. For hybrid agents (organoids, soft robots, and human–AI organisations) the design problem is to impose boundary conditions, constraining parts just enough to elicit target collective behaviour while preserving a non-empty intersection of viable policies (collective identity). When that intersection collapses, the system can fragment into competing subagents (Section 6).
@@
Future work should test these predictions by varying which layers are permitted to adapt (e.g. hardware-in-the-loop learning, morphologically adaptive substrates, and biologically integrated controllers) and measuring resulting degrees of freedom, free energy, and robustness. More broadly, this paper is a stepping stone toward a unified account of delegation, weakness/max-entropy heuristics, and multiscale agency in the larger Stack Theory programme [bennett2025thesis].
@@
To speculate about future research beyond experimental verification, Stack Theory is a lens that can be applied generally. For example, a number of software-first approaches to AI already approximate aspects of this delegation picture, including federated learning [nori2025federated], constructivist AGI programmes [thorisson2012, goertzel2014], and active inference [Friston2010, friston2013]. An expanded discussion (including AERA, Hyperon/ActPC-Chem, and links to scale-free active inference) is provided in the ESM (Supplementary Note S6). Stack Theory might also inform the design of ‘soft robots’ [man2019]. Soft robots are a proposed class of robots made of ‘soft’ materials that adapt. I have formalised adaptive materials as adaptation lower in the stack, which can be used as a formal foundation for soft robotics. Self-assembling nanoparticle networks [borghi2024_brainlikehw] are a step in this direction. They can be used to form reconfigurable logic gates [paroli2023_receptron], facilitating adaptation at very low levels of abstraction. Taken to its logical conclusion, future AI may more closely resemble artificial life [suzuki2006, ikegami2008, ikegami2007simulating]. Finally, Stack Theory may be extended to economic and organisational systems. For example, free-market trade could also be seen as a classic argument for delegation. As Hayek pointed out [hayek1945use], a great deal of relevant information is spread among a population, unavailable to a central planner. This motivates delegation of control, to shorten feedback loops. However, the present formalism treats resources implicitly via task correctness, so connecting these arguments to economic dynamics would require extending Stack Theory to explicit time-series [bennett2026a] and resource accumulation. Additional speculative extensions and applications (including organisational and economic systems, AI safety regulation, and temporal extensions of stacks) are collected in ESM Note S9.
In summary, I have applied Stack Theory to explore how delegation of control affects adaptability. The Law of the Stack shows how adaptability at higher levels of abstraction is constrained by adaptability at lower levels, and delegation is structurally necessary to minimise free energy. A practical design implication is that, when adaptability is desired, constraints should be weakened as far as possible without breaking the system’s intended function. Constraints that are not strictly necessary relative to the design goal should be removed, though what counts as “strictly necessary” is relative to design goals and value judgements (See ESM Supplementary Note S4.). Humans often prefer ‘additive’ solutions [adams2021], meaning if a problem can be solved by adding a rule or subtracting a rule, we tend to add a rule. However The Law of the Stack suggests a preference for subtractive solutions might be beneficial in many circumstances. The Law of the Stack might also be viewed as a formal justification of the classical principle of subsidiarity, often traced back to Aristotle, namely that higher levels of organization should coordinate and enable lower levels rather than replace functions they can adequately perform [aristotle_politics]. It points out a principle common to many systems, formalised in terms of Stack Theory and structurally tied to the FEP. Hence this paper lays a foundation for a broad range of future theoretical and empirical research, showing how systems of any kind might be made more robust and adaptable.
@@ Acknowledgements
I thank the anonymous peer reviewers for their detailed and helpful feedback. It made a substantial difference. I also thank Anna Ciaunica, Peter Watts, Sean Welsh, Ricard Solé, Karl Friston, Antonio Damasio, Kingson Man, Paolo Milani, Ben Goertzel, Kristinn R. Thórisson, Elija Perrier and Noel Hinton for their support and feedback on this paper.